Hello again,
I was just wondering about the answer to this problem 10-3w+1=9w+3-6w. When i did this problem, I got 8w=3. I was also hoping that you can give me the answer in steps as well as the check to see if it's right. thank you so much.

Answer by rfer(16322)

11-3w=3w+3
-6w=-8
w=-8/-6
w=1.333
------
proof
10-3*1.333+1=9*1.333+3-6*1.333
11-3.999=11.997+3-7.999
7=7
Slight error because of rounding